# Guaranteed Generation from Large Language Models

### v1

@@ Abstract
footnotetext: Work done during an internship at NAVER Labs Europe.footnotetext: Corresponding authors.
@@ 1 Introduction
In order to address these issues, we propose GUARD, [^1] The implementation of our framework, along with the code to replicate our experiments, will be made available soon. a simple framework for guaranteed generation that combines approximation of the ideal distribution $g$ into a proposal distribution $a^{\prime}$ with rejection sampling from the latter proposal. We establish a theorem (Theorem 2) about the relationship among the proposal $a^{\prime}$, the GUARD output $g^{\prime}$, and the ideal distribution $g$ through information geometry, by using the Pythagorean theorem for Information Projections (I-projections) [5]. As shown in Figure 1, this theorem states how it is possible to optimize both inference speed at sampling time (by improving the acceptance rate) and distributional closeness while strictly satisfying guarantees, by optimizing the approximation of the proposal $a^{\prime}$ towards $g$. We present several approaches for this approximation, stressing in particular the application to our situation of the Distributional Policy Gradient (DPG) technique [33, 13, 15, 16, 8] for training a parametric model to approximate an arbitrary unnormalized target distribution represented by an EBM.
@@
    Experiments with lexical and positive ending constraints, demonstrating that GUARD both significantly improves the acceptance rate and provides an excellent approximation to $g$; an analysis of the DPG and prompting approximation techniques, and the apparent limitations of prompting with respect to the diversity of $g$ in contrast to DPG; the warm-start combination of both techniques to speed up the approximation process.
@@ 2.1 Definition of the gold model
It is easy to see that $g$ is simply the distribution $a$ renormalized over the set of $y$’s that satisfy the constraint $b(y)=1$. We can characterize $g$ equivalently as follows (see App. A.1 for details):
@@ 2.2 From gold model to sampler
In order to associate a sampler to $g$, several approaches are possible. The first one consists of trying to find an autoregressive model (ARM) whose distribution is identical to $g$. Unfortunately, this is impossible in general, as we show below. Alternatively, an ARM $a^{\prime}$ could be considered in order to provide a “good enough” approximation to $g$, but it is far from obvious how to ensure that such a sampler will both satisfy the constraint $b$ and remain close to the gold model. A third approach is to use the ARM $a^{\prime}$ as a “proposal” inside some Monte-Carlo (MC) inference-time procedure. The framework that we propose, GUARD, is of this kind. It strictly guarantees the constraint and attempts to remain close to the gold model, while minimizing inference costs.
@@ 2.3 Limits of autoregressive models for filtering purposes
However, fine-tuning is not the only way to produce autoregressive models. Another approach would be as follows: we could, at each timestep $t$ of the generation of $y$, remove the token “bad” from the softmax vector $p(\cdot|y_{<t})$, and renormalize this vector to sum to 1. This would result in an ARM $a^{\prime}$, strictly satisfying the constraint, but would actually cause $a^{\prime}$ to have a distribution different from $g$, a phenomenon detailed in App. A.2.
So, obtaining an autoregressive model with distribution $g$ is not obvious, even in very simple cases as above, and we discuss in App. A.2 why this is a widespread problem in practice. As we mention there, we are only aware of very few special cases where the problem is solvable, but, as we will now discuss, we do have a clear negative result: the problem has no general solution.
@@
The following result was inspired by the pioneering work of Lin et al. [21] on the general limitations of autoregressive models. See App. A.3 for a precise statement and self-contained proof, adapted to the case of filtered models.
@@
There exists an autoregressive model $a$ and a binary predicate $b$ such that the gold model $g$ associated with $a$ and $b$ cannot be obtained by an autoregressive model $a^{\prime}$.
@@ 2.4 Claim: inference-time methods are needed
While it can be difficult or even impossible to exactly fit $g$ with an ARM $a^{\prime}$, it is still possible to approximate it with such a model and this can be done with various approaches such as (i) training $a^{\prime}$ through SFT or DPG [33], or else (ii) through prompting $a$ with an instruction or few-shot examples attempting to express the constraint $b$ in natural language. While the resulting model $a^{\prime}$ has the advantage that it can be used efficiently for generation, it typically does not strictly enforce $b$, our primary requirement, and we are not aware of general techniques that would allow an ARM $a^{\prime}$ to strictly enforce $b$ while simultaneously limiting the distortion of $a^{\prime}$ relative to $g$.
@@ 3 The GUARD framework and its properties
Inside the GUARD framework, we assume that $a^{\prime}$ is some autoregressive approximation to $g$. The GUARD sampler (Algorithm 1) is then an elementary form of rejection sampling. We assume that there exists at least one $y\in{\mathcal{Y}}$ s.t. $a^{\prime}(y)>0$ and $b(y)=1$. We then sample from $a^{\prime}$ until we find such a $y$, and return it. Such a procedure obviously guarantees that any sample $y$ satisfies the constraint, our fundamental requirement.
Algorithm 1 GUARD sampler
1:while True do
2:     $y\sim a^{\prime}$
3:     if $\textbf{if\ }b(y)=1\textbf{\ return\ }y$
We remark that in the special case where $a^{\prime}=a$, we are simply sampling from $a$ until we satisfy the constraint, and it is then easy to show that the sampler we obtain has exactly the same distribution as $g$ (see App. A.4). In this case, the GUARD sampler meets two criteria: (1) it strictly satisfies the constraint, and (2) its associated distribution is close to $g$ — in fact, it is equal to $g$. However, the third important criterion, efficiency, falls short when $a$ rarely satisfies the constraint $b$, which is equivalent to $\text{KL}(g||a)$ being large (see Eq. 7 in App. A.5).
By approximating $g$ with some $a^{\prime}$ that has a small $\text{KL}(g||a^{\prime})$ relatively to $\text{KL}(g||a)$ — in other words, $a^{\prime}$ is a better approximation to $g$ than $a$ — we can greatly improve the efficiency of GUARD, by sacrificing exact distributional match to $g$. Our core Theorem 2 below (see also the rightmost panel of Figure 1) characterizes this trade-off formally. Here, for a given $a^{\prime}$, $g^{\prime}=g^{\prime}_{a^{\prime}}$ denotes both the sampler resulting from Algorithm 1 and the associated distribution. $AR_{a^{\prime}}\doteq\mathbb{E}_{y\sim a^{\prime}}b(y)$ denotes the acceptance rate of the algorithm, that is, the probability that a sample from $a^{\prime}$ respects constraint $b$ and thus is accepted (see App. A.5).
@@
Proof sketch (see App. A.6 for details). The first equality results from the fact that $g^{\prime}$ is the I-projection of $a^{\prime}$ on ${\mathcal{C}}$ and from the Pythagorean identity for I-projections [5]. The second identity is obtained from a simple derivation showing that $\text{KL}(g^{\prime}||a^{\prime})=-\log Z^{\prime}=-\log AR_{a^{\prime}}$, where $Z^{\prime}$ is the partition function $Z^{\prime}\doteq\sum_{y\in{\mathcal{Y}}}a^{\prime}(y)\,b(y)$.
@@
This last observation motivates the need to minimize $\text{KL}(g||a^{\prime})$ to obtain a suitable $a^{\prime}$. We consider several methods to achieve this. (1) Prompting can be used to encourage the output to respect the given constraint — a technique we refer to as constraint-aware prompting (CAP). (2) Another approach consists in using Supervised Fine-Tuning (SFT). This involves sampling a large number of $y$’s from $a$, filtering them based on $b$ to create a dataset representing the $g$ distribution, and then fine-tuning $a$ on this dataset. However, when $AR_{a}$ is low, very few samples are retained, making this approach inefficient and difficult to use in practice, although it may still be useful for scientific analysis and comparisons. (3) The last method we consider is the Distributional Policy Gradient (DPG) technique for general EBMs [33, 16, 8] here applied to the specific case of filtered models. This method samples $y$ from a proposal $a^{\prime}$ initialized with $a$ and updates $a^{\prime}$ by performing gradient descent on $\text{KL}(g||a^{\prime})$. It is important to mention that one key difference between DPG and SFT is that DPG is adaptive; as the proposal $a^{\prime}$ approaches $g$, the frequency and quality of gradient updates increase, making it more efficient than SFT [13]. While the original DPG technique shares SFT’s initial low $AR_{a}$ during early training, we can improve this behavior by using constraint-aware prompting for the initial proposal — a method we refer to as “warm-start DPG” — to accept a rich amount of samples, thereby providing abundant gradient signals. Based on our experiments, this combination appears to be the most effective for guaranteed generation, and it is the one that we advocate. The full algorithm for warm-start DPG training in GUARD is provided in App. B as Algorithm 2.
@@
Simple rejection sampling as in Algorithm 1 might at first sight seem naive relative to some more sophisticated techniques, such as Markov Chain Monte Carlo (MCMC) approaches. One such technique, Independent Metropolis-Hastings (IMH) [38], can use an autoregressive model as its proposal, and construct a random walk that converges to $g$ in the limit, while still guaranteeing a strict respect of the constraint. On inspection, however, each individual move of this random walk implies — and therefore is at least as costly as — a full run of Algorithm 1. We report in App. F experiments showing that IMH converges very slowly, making it an impractical approach, and motivating our choice of Algorithm 1 combined with a focus on finding a good approximator $a^{\prime}$ to $g$. We also present in App. F some experiments based on Quasi-Rejection Sampling (QRS) [7] — a technique also converging to $g$ in the limit — which, while still being more costly than Algorithm 1, is shown to present certain advantages over IMH.
@@ 4 Experiments
In this section, we describe the experiments conducted to evaluate the GUARD framework on simple guaranteed generation scenarios. As discussed in the previous section, we compare three approximation methods within GUARD: CAP, SFT, and DPG. [^5] Training and evaluation were done using the disco toolkit: https://github.com/naver/disco. We have three main desiderata on GUARD’s outputs: (1) strict adherence to constraint $b$, (2) similarity to the gold distribution $g$, and (3) limited inference cost.666 In terms of computation costs, reducing the inference cost related to the acceptance rate $AR_{a^{\prime}}$ is our main focus, because this is the one involved in deployment. However, we also consider the training cost, in terms of the sampling budget required for training the approximation $a^{\prime}$, which is independent of the inference cost. Since $g^{\prime}$ is filtered by rejection sampling in GUARD and thus satisfies the first criterion by definition, our main metrics in the experiments are the distributional closeness $\text{KL}(g||g^{\prime})$ and the inference cost defined by the acceptance rate $AR_{a^{\prime}}$. Additional details about our metrics are provided in App. C.
We are considering two scenarios, text generation under a lexical constraint (Section 4.1) and story generation under a positive ending constraint for sentiment reversal (Section 4.2), where samples from the base LLM $a$ rarely respect the constraint $b$ (i.e., $\text{KL}(g||a)=-\log AR_{a}$ is high).777 If the acceptance rate of the base LLM $a$ is high enough, we can simply perform rejection sampling on proposal $a$ without any training. However, when violations of constraint $b$ are frequent, alternative methods are needed. This is the setting we explore in our experiments.
@@ 4.1 Unconditional generation with lexical constraints
Our first task is to generate a text under a lexical constraint, i.e., such that a specific keyword is included in the generation. The LLM generates a text $y$ of 30 tokens, starting with the $\langle bos\rangle$ token. The lexical constraint $b(y)$ then either rejects or accepts this generated text based on whether it contains the string of the keyword.
We use Gemma-2B [37] as our base LLM $a$. We define the constraint $b$ as the binary function that is equal to 1 if and only if the generated text $y$ contains the string “amazing”, following the experimental setup from Khalifa et al. [13]. This constraint has an acceptance rate $AR_{a}$ of approximately 0.0023 under the original distribution $a$ with ancestral sampling and a temperature of 1.
@@
Figure 3 depicts the results of the trade-off between $AR_{a^{\prime}}$ and $\text{KL}(g||g^{\prime})$ under a fixed sampling budget (800K). The point labeled as $\tilde{g}$ reflects the upper-bound performance associated to an ideal (but non-existent) sampler of the gold model, which would have no distributional difference with $g$ while achieving an acceptance rate of 1 (i.e., a 435$\times$ improvement with respect to $a$). Both DPG and SFT show significantly improved acceptance rates in comparison to $a$. However, DPG consistently demonstrates superior distributional similarity, as measured by $\text{KL}(g||g^{\prime})$. Interestingly, the CAP approach enforces a high acceptance rate without any training procedure, but it results in distributions that greatly diverge from the target $g$. Nevertheless, when DPG is initialized with $a(\cdot|\text{CAP})$, it leverages the high initial $AR_{a^{\prime}}$ to enable sufficient sampling. This, in turn, allows for a rapid reduction of $\text{KL}(g||g^{\prime})$. As a result, the figure shows that this warm-start DPG strictly Pareto-dominates all SFT points. This means that for any given point representing an SFT result, warm-start DPG offers a better acceptance rate and a lower KL divergence simultaneously. In summary, the GUARD algorithm increases the acceptance rate by almost 180$\times$ compared to the base LLM $a$ while maintaining minimal distributional distortion.
@@
As shown in Table 1, samples obtained from $g$ and DPG-based $g^{\prime}$ exhibit similar diversity scores. However, samples filtered through constraint-aware prompting show a significantly narrower range of diversity. This reduced diversity is also evident in the positional bias analysis in Figure 4, which studies the relative position of the string “amazing” in the generated text. Samples from $g$ and DPG-based $g^{\prime}$ show a relatively uniform distribution of “amazing” across token positions, whereas samples prompted to include the string “amazing” led to a heavy concentration of this word in early positions. Additional experimental details and analysis with other lexical constraints are provided in App. D, along with examples of generated samples.

### v2

@@ Abstract
footnotetext: Work done during an internship at NAVER Labs Europe.footnotetext: Correspondence to {minbeomkim@snu.ac.kr, marc.dymetman@gmail.com}
@@ 1 Introduction
In order to address these issues, we propose GUARD, [^1] The code is available at https://github.com/naver/guard. a simple framework for guaranteed generation that combines approximation of the ideal distribution $g$ into a proposal distribution $a^{\prime}$ with rejection sampling from the latter proposal. We establish a theorem (Theorem 2) about the relationship among the proposal $a^{\prime}$, the GUARD output $g^{\prime}$, and the ideal distribution $g$ through information geometry, by using the Pythagorean theorem for Information Projections (I-projections) [5]. As shown in Figure 1, this theorem states how it is possible to optimize both inference speed at sampling time (by improving the acceptance rate) and distributional closeness while strictly satisfying guarantees, by optimizing the approximation of the proposal $a^{\prime}$ towards $g$. We present several approaches for this approximation, stressing in particular the application to our situation of the Distributional Policy Gradient (DPG) technique [37, 13, 16, 17, 8] for training a parametric model to approximate an arbitrary unnormalized target distribution represented by an EBM.
@@
    Experiments with lexical and positive ending constraints, demonstrating that using GUARD leads to significantly improving the acceptance rate over rejection sampling from the base LLM $a$, while providing an excellent approximation to $g$; an analysis of the DPG and prompting approximation techniques, and the apparent limitations of prompting with respect to the diversity of $g$ in contrast to DPG; the warm-start combination of both techniques to speed up the approximation process.
@@ 2.1 Definition of the Gold Model
It is easy to see that $g$ is simply the distribution $a$ renormalized over the set of $y$’s that satisfy the constraint $b(y)=1$. We can characterize $g$ equivalently as follows (see App. B.1 for discussion):
@@ 2.2 From Gold Model to Sampler
In order to associate a sampler to $g$, several approaches are possible. The first one consists of trying to find an autoregressive model (ARM) whose distribution is identical to $g$. Unfortunately, this is impossible in general, as we show below. Alternatively, an ARM $a^{\prime}$ could be considered in order to provide a “good enough” approximation to $g$, but it is far from obvious how to ensure that such a sampler will both satisfy the constraint $b$ and remain close to the gold model. A third approach is
to use the ARM $a^{\prime}$ as a “proposal” inside some Monte-Carlo (MC) inference-time procedure. The framework that we propose, GUARD, is of this kind. It strictly guarantees the constraint and attempts to remain close to the gold model, while minimizing inference costs.
@@ 2.3 Limits of autoregressive models for filtering purposes
However, fine-tuning is not the only way to produce autoregressive models. Another approach would be as follows: we could, at each timestep $t$ of the generation of $y$, remove the token “bad” from the softmax vector $p(\cdot|y_{<t})$, and renormalize this vector to sum to 1. This would result in an ARM $a^{\prime}$, strictly satisfying the constraint, but would actually cause $a^{\prime}$ to have a distribution different from $g$, a phenomenon detailed in App. B.2.
So, obtaining an ARM with distribution $g$ is not obvious, even in very simple cases as above, and we discuss in App. B.2 why this is a widespread problem in practice. As we mention there, we are only aware of a few special cases where the problem is solvable, but, as we will now discuss, we do have a clear negative result: the problem has no general solution.
@@
We define a function from $V^{*}$ to $\mathbb{R}^{d}$ as polynomial-time computable (PTC) if and only if it can be computed in polynomial time relative to the length of its argument. A binary predicate is said to be PTC if it is a PTC function of its argument. An ARM $a$ is considered PTC if and only if the computation of the softmax vector at each step $t$ is a PTC function of the prefix up to that step.
The following result was inspired by the pioneering work of Lin et al. [23] on the general limitations of ARMs. See App. B.3 for a detailed statement and self-contained proof, adapted to the case of filtered models.
@@
Under the assumption $P\neq NP$, there exists a PTC ARM $a$ and a PTC binary predicate $b$ such that no PTC ARM $a^{\prime}$ has the same distribution as $g$.
@@
Interpretation: In practice, the PTC condition is satisfied by all standard ARM architectures, from Recurrent Neural Networks to Transformers of different flavors. Then, in essence, the theorem implies that it is in general impossible to fit $g$ with an ARM under such architectures.
@@ 2.4 Claim: inference-time methods are needed
While it can be difficult or even impossible to exactly fit $g$ with an ARM $a^{\prime}$, it is still possible to approximate it with such a model and this can be done with various approaches such as (i) training $a^{\prime}$ through Supervised Fine-Tuning (SFT) or DPG [37], or else (ii) through prompting $a$ with an instruction or few-shot examples attempting to express the constraint $b$ in natural language. While the resulting model $a^{\prime}$ has the advantage that it can be used efficiently for generation, it typically does not strictly enforce $b$, our primary requirement, and we are not aware of general techniques that would allow an ARM $a^{\prime}$ to strictly enforce $b$ while simultaneously limiting the distortion of $a^{\prime}$ relative to $g$.
@@ 3 The GUARD framework and its properties
Inside the GUARD framework, we assume that $a^{\prime}$ is some autoregressive approximation to $g$. The GUARD sampler is then an elementary form of rejection sampling which can be described by the following algorithm: sample $y$ from $a^{\prime}$ until $b(y)=1$, then return this $y$ (Algorithm 1). We assume that there exists at least one $y\in{\mathcal{Y}}$ s.t. $a^{\prime}(y)>0$ and $b(y)=1$. Such a procedure obviously guarantees that any sample $y$ satisfies the constraint, our fundamental requirement.
We remark that in the special case where $a^{\prime}=a$, we are simply sampling from $a$ until we satisfy the constraint, and it is then easy to show that the sampler we obtain has exactly the same distribution as $g$ (see App. B.4). In this case, the GUARD sampler meets two criteria: (1) it strictly satisfies the constraint, and (2) its associated distribution is close to $g$ — in fact, it is equal to $g$. However, the third important criterion, efficiency, falls short when $a$ rarely satisfies the constraint $b$, which is equivalent to $\text{KL}(g||a)$ being large (see Eq. 7 in App. B.5).
By approximating $g$ with some $a^{\prime}$ that has a small $\text{KL}(g||a^{\prime})$ relatively to $\text{KL}(g||a)$ — in other words, $a^{\prime}$ is a better approximation to $g$ than $a$ — we can greatly improve the efficiency of GUARD, by sacrificing exact distributional match to $g$. Our core Theorem 2 below (see also the rightmost panel of Figure 1) characterizes this trade-off formally. Here, for a given $a^{\prime}$, $g^{\prime}=g^{\prime}_{a^{\prime}}$ denotes both the sampler resulting from Algorithm 3 and the associated distribution. $AR_{a^{\prime}}\doteq\mathbb{E}_{y\sim a^{\prime}}b(y)$ denotes the acceptance rate of the algorithm, that is, the probability that a sample from $a^{\prime}$ respects constraint $b$ and thus is accepted (see App. B.5).
@@
Proof sketch (see App. B.6 for details). The first equality results from the fact that $g^{\prime}$ is the I-projection of $a^{\prime}$ on ${\mathcal{C}}$ and from the Pythagorean identity for I-projections [5]. The second identity is obtained from a simple derivation showing that $\text{KL}(g^{\prime}||a^{\prime})=-\log Z^{\prime}=-\log AR_{a^{\prime}}$, where $Z^{\prime}$ is the partition function $Z^{\prime}\doteq\sum_{y\in{\mathcal{Y}}}a^{\prime}(y)\,b(y)$.
@@
This last observation motivates the need to minimize $\text{KL}(g||a^{\prime})$ to obtain a suitable $a^{\prime}$. We consider several methods to achieve this. (1) Prompting can be used to encourage the output to respect the given constraint — a technique we refer to as constraint-aware prompting (CAP). (2) Another approach consists in using Supervised Fine-Tuning (SFT). This involves sampling a large number of $y$’s from $a$, filtering them based on $b$ to create a dataset representing the $g$ distribution, and then fine-tuning $a$ on this dataset. However, when $AR_{a}$ is low, very few samples are retained, making this approach inefficient and difficult to use in practice, although it may still be useful for scientific analysis and comparisons. (3) The last method we consider is the Distributional Policy Gradient (DPG) technique for general EBMs [37, 17, 8] here applied to the specific case of filtered models. This method samples $y$ from a proposal $a^{\prime}$ initialized with $a$ and updates $a^{\prime}$ by performing gradient descent on $\text{KL}(g||a^{\prime})$. It is important to mention that while both DPG and SFT seek to minimize $\text{KL}(g||a^{\prime})$ (see App. B.7 for more details), one key difference between them is that DPG is adaptive: as the proposal $a^{\prime}$ approaches $g$, the frequency and quality of gradient updates increase, making it more efficient than SFT [13]. While the original DPG technique shares SFT’s initial low $AR_{a}$ during early training, we can improve this behavior by using constraint-aware prompting for the initial proposal — a method we refer to as “warm-start DPG” — to accept a rich amount of samples, thereby providing abundant gradient signals. Based on our experiments, this combination appears to be the most effective for guaranteed generation, and it is the one that we advocate. The full algorithm for warm-start DPG training in GUARD is provided in App. C as Algorithm 2.
@@
Simple rejection sampling as in Algorithm 3 might at first sight seem naive relative to some more sophisticated techniques, such as Markov Chain Monte Carlo (MCMC) approaches. One such technique, Independent Metropolis-Hastings (IMH) [42], can use an autoregressive model as its proposal, and construct a random walk that converges to $g$ in the limit, while still guaranteeing a strict respect of the constraint. On inspection, however, each individual move of this random walk implies — and therefore is at least as costly as — a full run of Algorithm 3. We report in App. G experiments showing that IMH converges very slowly, making it an impractical approach, and motivating our choice of Algorithm 3 combined with a focus on finding a good approximator $a^{\prime}$ to $g$. We also present in App. G some experiments based on Quasi-Rejection Sampling (QRS) [7] — a technique also converging to $g$ in the limit — which, while still being more costly than Algorithm 3, is shown to present certain advantages over IMH.
@@ 4 Experiments
In this section, we describe the experiments conducted to evaluate the GUARD framework on simple guaranteed generation scenarios. As discussed in the previous section, we compare three approximation methods within GUARD: CAP, SFT, and DPG. [^5] Training and evaluation were done using the disco toolkit: https://github.com/naver/disco. We have three main desiderata on GUARD’s outputs: (1) strict adherence to constraint $b$, (2) similarity to the gold distribution $g$, and (3) limited inference cost.666 In terms of computation costs, reducing the inference cost related to the acceptance rate $AR_{a^{\prime}}$ is our main focus, because this is the one involved in deployment. However, we also consider the training cost, in terms of the sampling budget required for training the approximation $a^{\prime}$, which is independent of the inference cost. Since $g^{\prime}$ is filtered by rejection sampling in GUARD and thus satisfies the first criterion by definition, our main metrics in the experiments are the distributional closeness $\text{KL}(g||g^{\prime})$ and the inference cost defined by the acceptance rate $AR_{a^{\prime}}$. Additional details about our metrics are provided in App. D. [^7] Note that in this work our evaluation focus is in terms of the closeness of $g^{\prime}$ to the target gold model $g$ rather than in terms of the performance of $g^{\prime}$ on downstream tasks like the ones proposed in benchmarks such as CommonGen [22] — tasks that we see as directly influencing the design of the constraint $b$, and therefore the gold model $g$, which is then approximated with $g^{\prime}$.
We are considering two scenarios, text generation under a lexical constraint (Section 4.1) and story generation under a positive ending constraint for sentiment reversal (Section 4.2), where samples from the base LLM $a$ rarely respect the constraint $b$ (i.e., $\text{KL}(g||a)=-\log AR_{a}$ is high).888 If the acceptance rate of the base LLM $a$ is high enough, we can simply perform rejection sampling on proposal $a$ without any training. However, when violations of constraint $b$ are frequent, alternative methods are needed. This is the setting we explore in our experiments.
@@ 4.1 Unconditional generation with lexical constraints
Our first task is to generate a text under a lexical constraint, i.e., such that a specific keyword is included in the generation. The LLM generates a text $y$ of 30 tokens, starting with the $\langle bos\rangle$ token. The lexical constraint $b(y)$ then either rejects or accepts this generated text based on whether it contains the string of the keyword. We use Gemma-2B [41] as our base LLM $a$. We define the constraint $b$ as the binary function that is equal to 1 if and only if the generated text $y$ contains the string “amazing”, following the experimental setup from Khalifa et al. [13]. This constraint has an acceptance rate $AR_{a}$ of approximately 0.0023 (i.e., on average, one sample is accepted every 435 samples drawn) under the original distribution $a$ with ancestral sampling and a temperature of 1.
Empirical results
@@
Figure 4 depicts the results of the trade-off between $AR_{a^{\prime}}$ and $\text{KL}(g||g^{\prime})$ under a fixed sampling budget (800K). The point labeled as $\tilde{g}$ reflects the upper-bound performance associated to an ideal (but non-existent) sampler of the gold model, which would have no distributional difference with $g$ while achieving an acceptance rate of 1 (i.e., a 435$\times$ improvement with respect to $a$). Both DPG and SFT show significantly improved acceptance rates in comparison to $a$. However, DPG consistently demonstrates superior distributional similarity, as measured by $\text{KL}(g||g^{\prime})$. Interestingly, the CAP approach enforces a high acceptance rate without any training procedure, but it results in distributions that greatly diverge from the target $g$. Nevertheless, when DPG is initialized with $a(\cdot|\text{CAP})$, it leverages the high initial $AR_{a^{\prime}}$ to enable sufficient sampling. This, in turn, allows for a rapid reduction of $\text{KL}(g||g^{\prime})$. As a result, the figure shows that this warm-start DPG strictly Pareto-dominates all SFT points. This means that for any given point representing an SFT result, warm-start DPG offers a better acceptance rate and a lower KL divergence simultaneously. In summary, the GUARD algorithm increases the acceptance rate from 0.0023 to 0.416 (almost 180$\times$) compared to the base LLM $a$ while maintaining minimal distributional distortion.
@@
As shown in Table 1, samples obtained from $g$ and DPG-based $g^{\prime}$ exhibit similar diversity scores. However, samples filtered through constraint-aware prompting show a significantly narrower range of diversity. This reduced diversity is also evident in the positional bias analysis in Figure 2, which studies the relative position of the string “amazing” in the generated text. Samples from $g$ and DPG-based $g^{\prime}$ show a relatively uniform distribution of “amazing” across token positions, whereas samples prompted to include the string “amazing” led to a heavy concentration of this word in early positions. Additional experimental details and analysis with other lexical constraints are provided in App. E, along with examples of generated samples.